Lulabalula:

Lena is converting 2/25 to a percent. What should she multiply the numerator and denominator of the fraction by to find the percent?

Lulabalula:

@Imagine @supie @bigboybilly

imnotsmartlol:

why would someone ever do this in life

imnotsmartlol:

im gonna help you but why would anyone ever waste their time to convert things to 2/25

mxddi3:

Well, to turn this into a percent,we can see how many times 25 goes into 100. What is \[100\div25\] ?

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @imnotsmartlol im gonna help you but why would anyone ever waste their time to convert things to 2/25 \(\color{#0cbb34}{\text{End of Quote}}\) why would you be here if you wont help me I need help

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 Well, to turn this into a percent,we can see how many times 25 goes into 100. What is \[100\div25\] ? \(\color{#0cbb34}{\text{End of Quote}}\) 4

imnotsmartlol:

did u not read above it?

mxddi3:

yes, 4. so we are going to multiply them both by 4. \[\frac{ 2 }{ 25 }=\frac{ x }{ 100 }\] if we did 25x4=100 then what is 2x4 ?

bigboybilly:

A percent is a number out of 100, so you'd want to multiply the denominator to get 100. In this case, 25 would have to be multiplied by 4 to get 100. To keep the fractions equal to each other, you have to multiply the numerator by the same thing. 2*4=8. Therefore, your fraction would be 8/100 or 8%

imnotsmartlol:

woah smart

imnotsmartlol:

take that medal and wear it with honor mxddie

mxddi3:

thank you <3 please keep comments related to the post though!

imnotsmartlol:

¯\_(ツ)_/¯

imnotsmartlol:

oooo i need 2 more medals before 50

bigboybilly:

1 attachment